An astronaut uses a pendulum with a mass of 0.200 kg to measure the acceleration due to gravity on Planet X. He lifts the pendulum's mass a vertical height of 0.500 m and is able to determine that it gains 15.0 J of gravitational potential energy as it is lifted. Using this information, calculate the acceleration due to gravity (g) on Planet X. *




1
Expert's answer
2021-06-14T17:26:40-0400

By the definition of the gravitational potential energy we have:


"GPE=mgh,""g=\dfrac{GPE}{mh},""g=\dfrac{15.0\ J}{0.2\ kg\cdot0.5\ m}=150\ \dfrac{m}{s^2}."